Plot the following data, then find the slope and i
Answer #1

5. The plot is given by,

6 4 ? -5 10 15 20 25 Position x (m)

If the straight line is denoted by,

F(x) = mx + b

The slope of the straight line is given by, m = 0.214 N/m and the interception is b = 0.567 N.

-----------------------------------------------------------------------------------------------------------------------------------

6. If we compare the equation with the given one,

F(x) = \frac{x}{K} + \frac{a}{K}

we get the values of K and a.

K = \frac{1}{m} = 4.66 \hspace{2mm} m/N

a = Kb = 4.66 \times 0.567 = 2.64 \hspace{2mm} m

---------------------------------------------------------------------------------------------------------------------------------

7. No data found

------------------------------------------------------------------------------------------------------------------------------------

8. If we just plot the data it will look like,

It is clear that the plot is not linear. Look at the equation,

M = \frac{k}{f} +1

if we take a new variable as x = 1/f, then the equation would be linear, as,

M = kx +1

The new data set will be,

x = 1/f    -----------------    M       

1        13.1
0.67                                 9.1
0.5    7.0
0.25 4.0
0.2                                   3.3
0.143 2.6

We plot this set of data and get,

The slope of the curve is m = 12.26 m and the intercept is b = 0.88

Clearly the plot is linear now. From the equation,

M = kx +1

we can easily say that the slope is k = m = 12.26 m.